### v3

# Extremes of generalized inversions and descents on permutation groups

Abstract.
The classical inversion statistic on symmetric groups is the sum of all indicators $\textbf{1}\{\pi(i)>\pi(j)\}$ for a random permutation $\pi=(\pi(1),\ldots,\pi(n))$ and the pairs $(i,j)$ with $1\leq i<j\leq n$. The descent statistic counts all $i\in\{1,\ldots,n-1\}$ with $\pi(i)>\pi(i+1)$. The number of inversions can be generalized by restricting the indicators to pairs $(i,j)$ with $i<j$ and $|i-j|\leq d$ for some $d\in\{1,\ldots,n-1\}$. Likewise, the number of descents can be generalized by counting all $i\in\{1,\ldots,n-d\}$ with $\pi(i)>\pi(i+d)$. These generalized statistics can be further extended to the signed and even-signed permutation groups. The bandwidth index $d$ can be chosen in dependence of $n$, and the magnitude of $d$ is significant for asymptotic considerations. It is known that each of these statistics is asymptotically normal for suitable choices of $d$. In this paper we prove the bivariate asymptotic normality and determine the extreme value asymptotics of generalized inversions and descents.
Key words and phrases:
extreme values, CLT, permutation statistics, Coxeter group
2010 Mathematics Subject Classification:
Primary: 60G70, 05A16; Secondary: 20F55
Philip Dörr is supported by the DFG (314838170, GRK 2297, “MathCoRe”).
1. Introduction
The numbers of inversions and descents are two important quantities of permutations. For a permutation $\pi\negmedspace:\{1,\ldots,n\}\rightarrow\{1,\ldots,n\},$ an inversion is a pair $(i,j)$ with $i<j,$ but $\pi(i)>\pi(j)$. A descent is an index $i$ with $\pi(i)>\pi(i+1)$, i.e., descents correspond to adjacent inversions. We now equip the symmetric group $S_{n}$ with the discrete uniform probability measure induced by the point masses ($\mathbb{P}(\{\pi\})=1/n!$ $\forall\pi\in S_{n}$. In this context, we denote the random numbers of inversions and descents as $X_{\mathrm{inv}}$ and $X_{\mathrm{des}}$. These random variables can also be represented with help of i.i.d. standard uniform variables $Z_{1},Z_{2},\ldots,\sim U(0,1)$, namely,
(1) $X_{\mathrm{inv}}$ $=\sum_{1\leq i<j\leq n}\textbf{1}\{Z_{i}>Z_{j}\}\,,$
$X_{\mathrm{des}}$ $=\sum_{i=1}^{n-1}\textbf{1}\{Z_{i}>Z_{i+1}\}\,.$
In (1), we have to sum the indicators $\textbf{1}\{Z_{i}>Z_{j}\}$ over all pairs $(i,j)$ with $i<j$. A class of generalized inversion statistics can be constructed by restricting the sum in (1) to pairs $(i,j)$ with $1\leq j-i\leq d$, for some $d\in\{1,\ldots,n-1\}$. Writing $\mathfrak{N}_{n,d}:=\{(i,j)\in\{1,\ldots,n\}^{2}\mid 1<j-i\leq d\}$, we can say that
(2) $X_{\mathrm{inv}}^{(d)}:=\sum_{(i,j)\in\mathfrak{N}_{n,d}}\textbf{1}\{Z_{i}>Z_{j}\}$
counts the so-called $d$-inversions. The term generalized inversions is an umbrella term for all $d$-inversions. By choosing $d=n-1$ or $d=1,$ it is seen that this class includes common inversions and descents. Moreover, $d$-descents or generalized descents are given by
(3) $X_{\mathrm{des}}^{(d)}:=\sum_{i=1}^{n-d}\textbf{1}\{Z_{i}>Z_{i+d}\}\,.$
Generalized inversions were first introduced by de Mari & Shayman [5] to describe the Betti numbers of Hessenberg subvarieties in regular complex-valued matrices. The first stochastic considerations for the random variables $X_{\mathrm{inv}}^{(d)}$ are due to Bona [2] and Pike [11], who computed the mean and variance and proved a central limit theorem (CLT).
Symmetric groups belong to the class of finite irreducible Coxeter groups, which have been classified by [4]. The concept of inversions and descents can be transferred to these groups, see [1, Section 1.4]. Two other important subfamilies of the finite irreducible Coxeter groups are the signed permutation groups $B_{n}$ and the even-signed permutation groups $D_{n}$. The group $B_{n}$ consists of all maps $\pi\negthickspace:\{1,\ldots,n\}\longrightarrow\{1,\ldots,n\}\cup\{-1,\ldots,-n\}$ for which $|\pi|$ is a permutation. Such a map is called a signed permutation. The groups $D_{n}$ are the subgroups of $B_{n}$ containing all signed permutations with an even number of negative signs. The random number of inversions on these groups can be represented with help of i.i.d. variables $Z_{1},\ldots,Z_{n}\sim U(-1,1)$ as follows:
(4a) $X_{\mathrm{inv}}^{B}$ $=\sum_{1\leq i<j\leq n}\textbf{1}\{Z_{i}>Z_{j}\}+\sum_{1\leq i<j\leq n}\textbf{1}\{-Z_{i}>Z_{j}\}+\sum_{i=1}^{n}\textbf{1}\{Z_{i}<0\}\,,$
(4b) $X_{\mathrm{inv}}^{D}$ $=\sum_{1\leq i<j\leq n}\textbf{1}\{Z_{i}>Z_{j}\}+\sum_{1\leq i<j\leq n}\textbf{1}\{-Z_{i}>Z_{j}\}\,.$
We refer to the three families $S_{n},B_{n},$ and $D_{n}$ as the classical Weyl groups. Recently, Meier & Stump [10] extended the concept of generalized inversions and descents to the groups $B_{n}$ and $D_{n}$, again proving a CLT.
The investigation of extreme value asymptotics for common inversions and descents was initiated in [7, 6], so we aim to extend this knowledge to generalized inversions and descents. We aim to prove Gumbel attraction for both the stand-alone statistics $X_{\mathrm{inv}}^{(d_{1})},X_{\mathrm{des}}^{(d_{2})}$ and the joint statistic $(X_{\mathrm{inv}}^{(d_{1})},X_{\mathrm{des}}^{(d_{2})})^{\top}$. Here, $d_{1}$ and $d_{2}$ can be either fixed or dependent on $n$. We will investigate the impact of the choice of $d$ for these results.
This paper is structured as follows. Section 2 gives basic properties on generalized inversions and descents, and introduces Hájek projections that are used as independent sum approximations. Section 3 deduces the bivariate CLT and the extreme value theorems for generalized inversions and descents. Section 4 gives the proofs of two auxiliary lemmas determining the choices of $d$ for which the Hájek projection serves as a working approximation. We use typical Landau notation for positive sequences $a_{n},b_{n}$ as follows:
  • $a_{n}=O(b_{n})$ means that $\limsup_{n\rightarrow\infty}a_{n}/b_{n}<\infty$.
  • $a_{n}=o(b_{n})$ means that $\lim_{n\rightarrow\infty}a_{n}/b_{n}=0$. This is also written as $a_{n}\ll b_{n}$ or $b_{n}\gg a_{n}$.
  • $a_{n}=\Theta(b_{n})$ means that $a_{n}$ and $b_{n}$ have the same order of magnitude, i.e., both $a_{n}=O(b_{n})$ and $b_{n}=O(a_{n})$ hold.
  • $a_{n}=b_{n}+o_{\mathbb{P}}(1)$ means that $a_{n},b_{n}$ are sequences of random variables with $a_{n}-b_{n}\overset{\mathbb{P}}{\longrightarrow}0$.
2. Basic definitions
Definition 2.1.
Let $S_{n}$ be a symmetric group and let $\pi\in S_{n}$. For $d\in\{1,\ldots,n-1\}$, $d$-inversions are all pairs $(i,j)$ in $\mathfrak{N}_{n,d}$ with $\pi(i)>\pi(j)$. In this sense, common descents equal $1$-inversions, and common inversions equal $(n-1)$-inversions. Moreover, $d$-descents are all numbers $i\in\{1,\ldots,n-d\}$ with $\pi(i)>\pi(i+d)$. We write $X_{\mathrm{inv}}^{(d)}$ for the random number of $d$-inversions and $X_{\mathrm{des}}^{(d)}$ for the number of $d$-descents. Probabilistic representations are given in (2), (3).  [^1] In the literature, there are different terminologies, e.g., in [2, 11], $d$-inversions are called $d$-descents. However, we use the terms and notation provided in [10].
Remark 2.2.
Obviously, every number $k\in\{1,\ldots,n\}$ can appear in at most $2d$ $d$-inversions. This bound is redundant if $d>n/2$. In fact, it is an important case distinction whether $d\leq n/2$ or $d>n/2$, e.g., when calculating the mean and variance of $X_{\mathrm{inv}}^{(d)}$ and $X_{\mathrm{des}}^{(d)}$. In the case of $d\leq n/2,$ we can split $\{1,\ldots,n\}$ into the regions
$K_{1}:=\{1,\ldots,d\},\qquad K_{2}:=\{d+1,\ldots,n-d\},\qquad K_{3}:=\{n-d+1,\ldots,n\}.$
For each $k\in K_{2},$ all larger indices $k+1,\ldots,k+d$ and all smaller indices $k-1,\ldots,k-d$ allow to form $d$-inversions. For any $k\notin K_{2},$ there are less than $d$ indices available in one direction. If $k\in K_{1}$, then $k-1$ smaller indices are available for $d$-inversions, which is less than $d$ numbers. If $k>n-d$ is large, then there are only $n-k<d$ larger indices available.
On the contrary, if $d>n/2,$ then $n-d<d$ and the above partition into three regions is now written as
$K_{1}:=\{1,\ldots,n-d\},\qquad K_{2}:=\{n-d+1,\ldots,d\},\qquad K_{3}:=\{d+1,\ldots,n\}.$
Now, if $k\in K_{2},$ then there are only $k-1$ smaller indices and $n-k$ larger indices available to form $d$-inversions.
The mean and variance of $X_{\mathrm{inv}}^{(d)}$ have been extensively computed by Pike [11]. It is easy to verify that the special cases $d=1$ and $d=n-1$ are consistent with [8, Corollaries 3.2 and 4.2]. The proof of [11, Theorem 1] is reviewed in [10, Theorem A.1], where the variance of $d$-descents is provided as well.
Theorem 2.3 (see [11], Theorem 1 and [10], Theorem A.1).
For all $d=1,\ldots,n-1,$ it holds that
$\mathbb{E}\left(X_{\mathrm{inv}}^{(d)}\right)$ $=\frac{2nd-d^{2}-d}{4},$ $\mathbb{E}\left(X_{\mathrm{des}}^{(d)}\right)$ $=\frac{n-d}{2}.$
Moreover, if $d\leq n/2,$ then
$\mathrm{Var}\left(X_{\mathrm{inv}}^{(d)}\right)$ $=\frac{6nd+4d^{3}+3d^{2}-d}{72},$ $\mathrm{Var}\left(X_{\mathrm{des}}^{(d)}\right)$ $=\frac{n+d}{12}.$
If $d>n/2,$ then
$\mathrm{Var}\left(X_{\mathrm{inv}}^{(d)}\right)=-\frac{1}{6}d^{3}+\left(\frac{1}{3}n-\frac{7}{24}\right)d^{2}-\left(\frac{1}{6}n^{2}-\frac{5}{12}n+\frac{1}{8}\right)d+\frac{1}{36}n^{3}-\frac{1}{12}n^{2}+\frac{1}{18}n\,$
and $\mathrm{Var}\left(X_{\mathrm{des}}^{(d)}\right)=(n-d)/4$.
Meier & Stump [10] introduced an extension of generalized inversions and descents from symmetric groups to the other classical Weyl groups $B_{n}$ and $D_{n}$. This extension is based on the root poset of a classical Weyl group. We refer to [10, Section 2] for the details. On the symmetric group $S_{n}$, the ordered pairs of indices $(i,j)$ correspond to the positive roots $[ij]:=e_{i}-e_{j}$, where $e_{i},e_{j}$ are unit vectors in $\mathbb{R}^{n}$, and the height of $[ij]$ within the root poset is $\text{ht}([ij])=j-i$.
On the signed permutation group $B_{n},$ we also have to consider the positive roots $[\widetilde{ij}]:=e_{i}+e_{j}$ and $[i]:=e_{i}$ for $1\leq i<j\leq n$. The heights of these additional roots are $\text{ht}([\widetilde{ij}])=i+j$ and $\text{ht}([i])=i$. The root $[\widetilde{ij}]$ corresponds to the indicator $\textbf{1}\{-\pi(i)>\pi(j)\}$ appearing in  (4a), while $[i]$ corresponds to $\textbf{1}\{\pi(i)<0\}$. See [10, Example 2.2] for an illustration of the root poset of $B_{n}$. On the even-signed permutation group $D_{n},$ the roots $[i]$ are disregarded, and $[\widetilde{ij}]$ has height $i+j-2$.
Definition 2.4.
For any classical Weyl group, $d$-inversions are determined by roots of height at most $d$, and $d$-descents are determined by roots of height exactly $d$, see [10, Definition 2.4]. For symmetric groups, this coincides with Definition 2.7. In addition to $\mathfrak{N}_{n,d}$, we introduce
$\widetilde{\mathfrak{N}}_{n,d}:=\{(i,j)\in\{1,\ldots,n\}^{2}\mid i<j,j+i\leq d\}.$
Then, on the signed and even-signed permutation groups, $X_{\mathrm{inv}}^{(d)}$ and $X_{\mathrm{des}}^{(d)}$ can be expressed as follows:
(5a) $X_{\mathrm{inv}}^{B,(d)}$ $=\sum_{(i,j)\in\mathfrak{N}_{n,d}}\textbf{1}\{Z_{i}>Z_{j}\}+\sum_{(i,j)\in\widetilde{\mathfrak{N}}_{n,d}}\textbf{1}\{-Z_{i}>Z_{j}\}+\sum_{i=1}^{d}\textbf{1}\{Z_{i}<0\}\,,$
(5b) $X_{\mathrm{inv}}^{D,(d)}$ $=\sum_{(i,j)\in\mathfrak{N}_{n,d}}\textbf{1}\{Z_{i}>Z_{j}\}+\sum_{(i,j)\in\widetilde{\mathfrak{N}}_{n,d+2}}\textbf{1}\{-Z_{i}>Z_{j}\}\,,$
(5c) $X_{\mathrm{des}}^{B,(d)}$ $=\sum_{i=1}^{n-d}\textbf{1}\{Z_{i}>Z_{i+d}\}+\sum_{i=1}^{\lceil d/2\rceil-1}\textbf{1}\{-Z_{i}>Z_{d-i}\}+\textbf{1}\{\pi(i)>d\}\,,$
(5d) $X_{\mathrm{des}}^{D,(d)}$ $=\sum_{i=1}^{n-d}\textbf{1}\{Z_{i}>Z_{i+d}\}+\sum_{i=1}^{\lceil d/2\rceil}\textbf{1}\{-Z_{i}>Z_{d+2-i}\}\,.$
Remark 2.5.
The largest possible choice of $d$ equals the total height of the root poset, namely, $d_{\max}-1$, where $d_{\max}$ denotes the largest degree of the underlying classical Weyl group. In particular, $d_{\max}=n$ for $S_{n},$ $d_{\max}=2n$ for $B_{n}$ and $d_{\max}=2n-2$ for $D_{n}$. To precisely compute the variance of $X_{\mathrm{inv}}^{(d)}$ and $X_{\mathrm{des}}^{(d)}$ on the groups $B_{n}$ and $D_{n},$ one needs to distinguish eight cases, as seen in [10, Theorems A.4 and A.13]. However, many of these cases give the same asymptotic quantification, which can be stated as follows:
Lemma 2.6 (cf. [10], Theorems A.4 and A.13).
For the generalized inversions and descents on both the groups $B_{n}$ and $D_{n},$ it holds that
$\mathrm{Var}\left(X_{\mathrm{inv}}^{(d)}\right)$ =$=\begin{cases}\frac{1}{36}d^{3}+\frac{1}{12}nd+O(d^{2}),&d\leq n/2\\ \frac{1}{36}d^{3}+O(d^{2}),&n/2\leq d<n\\ -\frac{1}{12}d^{3}+\frac{1}{3}nd^{2}-\frac{1}{3}n^{2}d+\frac{1}{9}n^{3}+O(d^{2}),&d\geq n\end{cases},$
$\mathrm{Var}\left(X_{\mathrm{des}}^{(d)}\right)$ =$=\begin{cases}\frac{1}{24}d+\frac{1}{12}n+O(1),&d<n\\ -\frac{1}{8}d+\frac{1}{4}n+O(1),&d\geq n\end{cases}.$
An evident issue seen in (2), (5a), and (5b) is the mutual dependence of the indicators $\textbf{1}\{Z_{i}>Z_{j}\}$ and $\textbf{1}\{-Z_{i}>Z_{j}\}$. Therefore, we aim to approximate $X_{\mathrm{inv}}^{(d)}$ with a sum of independent variables. Hájek projections are a common tool for this.
Definition 2.7.
Let $Z_{1},\ldots,Z_{n}$ be independent random variables, and let $X$ be another random variable. Then, the Hájek projection of $X$ with respect to $Z_{1},\ldots,Z_{n}$ is given by
$\hat{X}:=\sum_{k=1}^{n}\mathbb{E}(X\mid Z_{k})-(n-1)\mathbb{E}(X).$
As every $\mathbb{E}(X\mid Z_{k})$ is a measurable function only in $Z_{k}$, the Hájek projection is a sum of independent random variables. If $(X_{n})_{n\in\mathbb{N}}$ is a sequence of random variables with each $X_{n}$ depending on $Z_{1},\ldots,Z_{n}$, then the Hájek projections $\hat{X}_{n}$ should give a sufficiently accurate approximation to $X_{n}$. This can be ensured by the following criterion.
Theorem 2.8 (cf. [13], Theorem 11.2).
Consider a sequence $(X_{n})_{n\geq 1}$ of random variables and their associated Hájek projections $(\hat{X}_{n})_{n\geq 1}$. If $\mathrm{Var}(\hat{X}_{n})\sim\mathrm{Var}(X_{n})$ as $n\rightarrow\infty,$ then
$\frac{X_{n}-\mathbb{E}(X_{n})}{\mathrm{Var}(X_{n})^{1/2}}=\frac{\hat{X}_{n}-\mathbb{E}(\hat{X}_{n})}{\mathrm{Var}(\hat{X}_{n})^{1/2}}+o_{\mathbb{P}}(1).$
3. Asymptotic results
Asymptotic normality
For the univariate statistics $X_{\mathrm{inv}}^{(d)}$ and $X_{\mathrm{des}}^{(d)}$, the asymptotic normality has been proven in [10]. For the joint statistic $\left(X_{\mathrm{inv}}^{(d)},X_{\mathrm{des}}^{(d)}\right)^{\top}$, the simplest case is when $d$ remains fixed. Then, both $X_{\mathrm{inv}}^{(d)}$ and $X_{\mathrm{des}}^{(d)}$ are $m$-dependent $(m=d)$ and it is not necessary to apply the Hájek approximation. From the CLT for $m$-dependent random vectors, it follows that:
Theorem 3.1.
For any two fixed numbers $d_{1},d_{2},$ the joint distribution $\left(X_{\mathrm{inv}}^{(d_{1})},X_{\mathrm{des}}^{(d_{2})}\right)^{\top}$ satisfies the CLT.
An even stronger statement than the CLT is the uniform Gaussian approximation over all hyperrectangles for $m$-dependent random vectors given in [3, Theorem 2]. It also applies for random vectors with a sparse dependency graph. Let $\left(X_{j}^{(n)}\right)_{j=1,\ldots,n}$ be a triangular array of $p$-dimensional random vectors (which, without restriction, are centered) and let
$X^{(n)}$ $:=\sum_{t=1}^{n}X_{t}^{(n)},$ $\Sigma^{(n)}$ $:=\mathrm{Var}(X^{(n)})\,,$ $\mathcal{N}$ $\sim\mathrm{N}_{p}(0,\Sigma^{(n)})\,.$
Let $\mathcal{A}^{\mathrm{re}}$ be the system of all $p$-dimensional hyperrectangles, including infinite bounds, i.e., $\mathcal{A}^{\mathrm{re}}:=\bigl{\{}\{\textbf{w}\in\mathbb{R}^{p}\negmedspace:\textbf{a}\leq\textbf{w}\leq\textbf{b}\}\mid\textbf{a},\textbf{b}\in[-\infty,\infty]^{p}\bigr{\}}\,.$ Let
$r_{n}(\mathcal{A}^{\mathrm{re}}):=\sup_{A\in\mathcal{A}^{\mathrm{re}}}|\mathbb{P}(X^{(n)}\in A)-\mathbb{P}(\mathcal{N}\in A)|\,.$
In particular, $r_{n}(\mathcal{A}^{\mathrm{re}})$ is an upper bound for
$r_{n}(\mathcal{A}^{\mathrm{CLT}}):=\sup_{u\in\mathbb{R}^{p}}|\mathbb{P}(X^{(n)}\leq u)-\mathbb{P}(\mathcal{N}_{n}\leq u)|\,,$
which is relevant for the CLT, and for
$r_{n}(\mathcal{A}^{\mathrm{ext}}):=\sup_{u\in\mathbb{R}^{p}}|\mathbb{P}(X^{(n)}>u)-\mathbb{P}(\mathcal{N}_{n}>u)|\,,$
which is relevant for the asymptotics of extreme values. Now, let $G_{n}$ be a dependency graph for $X_{1}^{(n)},\ldots,X_{n}^{(n)},$ which consists of all edges $(i,j)$ for which $X_{i}^{(n)}$ and $X_{j}^{(n)}$ are dependent. Let $\Delta_{n}$ be the maximum degree of $G_{n}$ and let $\Delta_{n}^{*}$ be the maximum degree of the 2-reachability graph of $G_{n}$. If the graphs $G_{n}$ are not too dense as $n\rightarrow\infty,$ then $r_{n}(\mathcal{A}^{\mathrm{re}})$ can be bounded as follows:
Theorem 3.2 (cf. [3], Theorem 2).
Let $(X_{j}^{(n)})_{j=1,\ldots,n},$ $r_{n}(\mathcal{A}^{\mathrm{re}}),$ $\Delta_{n},\Delta_{n}^{*}$ be as above. Under some regularity conditions which are satisfied by bounded variables, it holds that
$r_{n}(\mathcal{A}^{\mathrm{re}})=O\left(\frac{(\Delta_{n}\Delta_{n}^{*})^{1/3}\log(p)^{7/6}}{n^{1/6}}\right).$
In particular, if $(X_{j}^{(n)})_{j=1,\ldots,n}$ is $m$-dependent for some global constant $m\in\mathbb{N},$ then $r_{n}(\mathcal{A}^{\mathrm{re}})=O\bigl{(}n^{-1/6}\log(p)^{7/6}\bigr{)}$.
This theorem was stated by Chang et al. for high dimensions, but it also works with fixed $p$ by artificially repeating the components of a random vector in fixed dimension.
From now, we consider the univariate statistics $X_{\mathrm{inv}}^{(d_{1})},X_{\mathrm{des}}^{(d_{2})}$ and the joint statistic $\left(X_{\mathrm{inv}}^{(d_{1})},X_{\mathrm{des}}^{(d_{2})}\right)^{\top}$ for two sequences $d_{1}=d_{1}(n),$ $d_{2}=d_{2}(n)$ with $d_{1}(n),d_{2}(n)\leq d_{\max}-1$ $\forall n\in\mathbb{N}$. Both $X_{\mathrm{inv}}^{(d_{1})}$ and $X_{\mathrm{des}}^{(d_{2})}$ are based on a sequence of classical Weyl groups $(W_{n})_{n\in\mathbb{N}},$ where each $W_{n}$ is one of $S_{n},B_{n},$ or $D_{n}$.
If $d_{2}=d_{2}(n)$ grows as $n\rightarrow\infty$, then there exists no constant $m\in\mathbb{N}$ for which all $X_{\mathrm{des}}^{(d_{2})}$ are $m$-dependent. However, the dependency structure of $X_{\mathrm{des}}^{(d_{2})}$ is still sparse. Recall that $X_{\mathrm{inv}}^{(d_{1})}$ and $X_{\mathrm{des}}^{(d_{2})}$ are constructed from i.i.d. random variables $Z_{1},\ldots,Z_{n}$. According to (3), (5c), (5d), we can represent $X_{\mathrm{des}}^{(d_{2})}$ as a sum of indicator variables, each of which depends on at most three others. Therefore, the first-order and second-order maximum degrees $\Delta_{n},\Delta_{n}^{*}$ are bounded in the way of $\Delta_{n}\leq 3$ and $\Delta_{n}^{*}\leq 9$. In conclusion, there are no issues with applying Theorem 3.2 if $d_{2}$ grows.
If $d_{1}=d_{1}(n)$ grows as well, then the maximum degrees $\Delta_{n},\Delta_{n}^{*}$ of $\left(X_{\mathrm{inv}}^{(d_{1})},X_{\mathrm{des}}^{(d_{2})}\right)^{\top}$ are bounded in the way of $\Delta_{n}\leq 4d_{1}$ and $\Delta_{n}^{*}\leq 8d_{1}$. Moreover, we have to take into account that by (2), (5a), (5b), $X_{\mathrm{inv}}^{(d_{1})}$ is based on $\Theta(nd_{1})$ summands, so we have to replace $n$ with $nd$ in Theorem 3.2. Therefore, Theorem 3.2 gives an $o(1)$ bound of $r_{n}(\mathcal{A}^{\mathrm{re}})$ if and only if $(d_{1})^{2/3}=o((nd_{1})^{1/6})\Longleftrightarrow d_{1}=o(n^{1/3})$. For any faster growth rate of $d_{1},$ the dependency structure of $\left(X_{\mathrm{inv}}^{(d_{1})},X_{\mathrm{des}}^{(d_{2})}\right)^{\top}$ is too complex to apply Theorem 3.2. In that case, we have to replace $X_{\mathrm{inv}}^{(d_{1})}$ with the Hájek projection $\hat{X}_{\mathrm{inv}}^{(d_{1})}$. The growth rate of $d_{1}$ determines whether the condition in Theorem 2.8 is fulfilled or not. To classify the validity of this condition, we distinguish the short case $d\leq d_{\max}/2$ and the long case $d>d_{\max}/2$.
Lemma 3.3.
Given the short case, we have for symmetric groups that
$\mathrm{Var}\left(\hat{X}_{\mathrm{inv}}^{(d)}\right)=\frac{1}{72}\left(4d^{3}+6d^{2}+2d\right)\,,$
and we have $\mathrm{Var}\left(\hat{X}_{\mathrm{inv}}^{(d)}\right)=d^{3}/36+O(d^{2})$ for the groups $B_{n}$ and $D_{n}$. So, if $d_{\max}/2>d\gg\sqrt{n},$ then $\mathrm{Var}\left(\hat{X}_{\mathrm{inv}}^{(d)}\right)\sim\mathrm{Var}\left(X_{\mathrm{inv}}^{(d)}\right)$.
Lemma 3.4.
Given the long case, $\mathrm{Var}\left(\hat{X}_{\mathrm{inv}}^{(d)}\right)\sim\mathrm{Var}\left(X_{\mathrm{inv}}^{(d)}\right)$ holds on symmetric groups if and only if $d\sim d_{\max}$. On the other classical Weyl groups, $\mathrm{Var}\left(\hat{X}_{\mathrm{inv}}^{(d)}\right)\sim\mathrm{Var}\left(X_{\mathrm{inv}}^{(d)}\right)$ holds if and only if $d\sim d_{\max}$ or $d\sim d_{\max}/2$.
From Theorem 3.2 and Slutsky’s lemma, we can conclude:
Corollary 3.5.
$\left(X_{\mathrm{inv}}^{(d_{1})},X_{\mathrm{des}}^{(d_{2})}\right)^{\top}$ satisfies the CLT if $d_{1}=o(n^{1/3})$ or if $d_{1}$ satisfies the conditions of either Lemma 3.3 or Lemma 3.4.
Extreme values of the univariate statistics $X_{\mathrm{inv}}^{(d)}$ and $X_{\mathrm{des}}^{(d)}$
We now postulate the univariate and bivariate extreme value limit theorems (EVLTs) for $X_{\mathrm{inv}}^{(d)}$ and $X_{\mathrm{des}}^{(d)}$ with the help of the methods used in [6]. The proof of [6, Theorem 4.1] consists of two parts. The first part is to show the Gumbel attraction of the statistic that replaces $X_{\mathrm{inv}}^{(d)}$ with $\hat{X}_{\mathrm{inv}}^{(d)}$. This is done by an application of [3, Theorem 2], providing Gaussian approximation for $m$-dependent random vectors. The second part is to show that the error resulting from this replacement vanishes in probability. Recall (e.g., by [9], Theorem 1.5.3]) that by use of
$\alpha_{n}$ $=\frac{1}{\sqrt{2\log n}}\,,$ $\beta_{n}$ $=\frac{1}{\alpha_{n}}-\frac{1}{2}\alpha_{n}\bigl{(}\log\log n+\log(4\pi)\bigr{)}\,,$
the maximum of $n$ i.i.d. variables $N_{1},\ldots,N_{n}\sim\mathrm{N}(0,1)$ is attracted to the Gumbel distribution, that is, $\forall x\in\mathbb{R}$:
$\mathbb{P}\left(\max_{i=1,\ldots,n}N_{i}\leq\alpha_{n}x+\beta_{n}\right)\longrightarrow\exp(-\exp(-x))=:\Lambda(x).$
Likewise, the two-dimensional standard normal distribution is attracted to the two-dimensional Gumbel distribution with independent marginals. Taking $\boldsymbol{\upalpha}_{n}:=(\alpha_{n},\alpha_{n}),$ $\boldsymbol{\upbeta}_{n}:=(\beta_{n},\beta_{n})\in\mathbb{R}^{2}$ and writing ”*” for component-wise multiplication, we have by [12, Theorem 3] that for i.i.d. bivariate standard normal $\mathbf{N}_{1},\ldots,\mathbf{N}_{n}$ and $\mathbf{x}=(x_{1},x_{2})\in\mathbb{R}^{2}$:
$\mathbb{P}\left(\max_{i=1,\ldots,n}\mathbf{N}_{i}\leq\boldsymbol{\upalpha}_{n}*\mathbf{x}+\boldsymbol{\upbeta}_{n}\right)\longrightarrow\exp\bigl{(}-\exp(-x_{1})-\exp(-x_{2})\bigr{)}=:\Lambda_{2}(x).$
Since the random numbers of generalized inversions and descents are discrete distributions, we are interested in the extremes of a triangular array $(X_{n1},\ldots,X_{nk_{n}}),$ where for each $n\in\mathbb{N},$ the block $X_{n1},\ldots,X_{nk_{n}}$ consists of i.i.d. samples taken from the $n$-th symmetric group $S_{n}$. Each of the following EVLTs imposes an upper bound on the number of samples $k_{n}$.
For a univariate triangular array consisting of generalized descents, it is not necessary to use the Hájek projection. It has already been argued in [6, Remark 4.2] that a subexponential bound of $k_{n}$ can be obtained if the Hájek projection is not needed. So, in the univariate EVLT for generalized descents, we can use the bound of $k_{n}=\exp(o(n^{1/7}))$ stated in [6, Remark 4.2]:
Theorem 3.6.
Let $(X_{nj})_{j=1,\ldots,k_{n}}$ be a row-wise i.i.d. triangular array with $X_{n1}\overset{\mathcal{D}}{=}X_{\mathrm{des}}^{(d)}$ for any sequence $d=d(n)$ with $1\leq d\leq d_{\max}-1,$ and let $M_{n}:=\max\{X_{n1},\ldots,X_{nk_{n}}\}$. Let $a_{n}:=\sigma\bigl{(}X_{\mathrm{des}}^{(d)}\bigr{)}\alpha_{k_{n}}$ and let $b_{n}:=\sigma\bigl{(}X_{\mathrm{des}}^{(d)}\bigr{)}\beta_{k_{n}}+\mu\bigl{(}X_{\mathrm{inv}}^{(d)}\bigr{)}.$ If $k_{n}=\exp(o(n^{1/7})),$ then
$\forall x\in\mathbb{R}\negmedspace:\quad\mathbb{P}(M_{n}\leq a_{n}x+b_{n})\longrightarrow\Lambda(x)\,.$
A similar statement applies to $X_{\mathrm{inv}}^{(d)}$ if $d$ grows slow enough to permit the application of Theorem 3.2.
Theorem 3.7.
Let $(X_{n1},\ldots,X_{nk_{n}})$ be a row-wise i.i.d. triangular array with $X_{n1}\overset{\mathcal{D}}{=}X_{\mathrm{inv}}^{(d)}$, where $d=d(n)=o(n^{1/3})$. Let $M_{n},a_{n},b_{n}$ be given in analogy to Theorem 3.6. If $k_{n}=\exp\bigl{(}o(n^{1/7}/d^{3/7})\bigr{)},$ then
$\forall x\in\mathbb{R}\negmedspace:\quad\mathbb{P}(M_{n}\leq a_{n}x+b_{n})\longrightarrow\Lambda(x)\,.$
Proof.
According to the above considerations, the maximum degrees $\Delta_{n},\Delta_{n}^{*}$ in the dependency graph of the representations (2), (5a), and (5b) are bounded in the way of $\Delta_{n}\leq 4d$ and $\Delta_{n}^{*}\leq 8d$. Since these representations are based on $\Theta(nd)$ summands, we need to replace $n$ with $nd$ when applying Theorem 3.2. An application of Theorem 3.2 with $\max\{n,k_{n}\}$ i.i.d. iterations of $X_{\mathrm{inv}}^{(d)}$ yields
$|\mathbb{P}(M_{n}\leq a_{n}x+b_{n})-\mathbb{P}(\mathcal{M}_{n}\leq\alpha_{n}x+\beta_{n})|=O(n^{-1/6}d^{1/2})\log(k_{n})^{7/6}=o(1)\,,$
where $\mathcal{M}_{n}$ is the maximum of $n$ i.i.d. copies of the standard normal distribution. The claim follows. ∎
For any other growth rate of $d,$ we can state an EVLT only for the cases covered by Lemmas 3.3 and 3.4, and due to the use of Hájek’s projection, we can only impose a strongly reduced asymptotic bound for $k_{n}$.
Theorem 3.8.
Let $(X_{n1},\ldots,X_{nk_{n}})$ be a row-wise i.i.d. triangular array with $X_{n1}\overset{\mathcal{D}}{=}X_{\mathrm{inv}}^{(d)}$, where $d=d(n)$ satisfies the conditions in either Lemma 3.3 or 3.4. Assume $k_{n}\log(k_{n})=o(n^{-1}d^{2})$ in the short case and $k_{n}\log(k_{n})=o\left(\min\left\{n,\displaystyle{\frac{n}{n-d}}\right\}\right)$ in the long case. Let $M_{n}:=\max\{X_{n1},\ldots,X_{nk_{n}}\}$ be the row-wise maximum. Then,
$\forall x\in\mathbb{R}\negthickspace:\quad\mathbb{P}(M_{n}\leq a_{n}x+b_{n})\longrightarrow\Lambda(x)\,,$
with $a_{n},b_{n}$ given in analogy to Theorem 3.6.
Proof.
The conditions on $d$ ensure that $1-\mathrm{Var}(X_{\mathrm{inv}}^{(d)})/\mathrm{Var}(\hat{X}_{\mathrm{inv}}^{(d)})=o(1),$ but the rate of convergence determines the bound for $k_{n}$ by means of [6, Eq. (9), (10)]. We compute this rate for symmetric groups, since the same conclusions can be obtained on the other classical Weyl groups. In the short case, we have
$1-\frac{\mathrm{Var}(X_{\mathrm{inv}}^{(d)})}{\mathrm{Var}(\hat{X}_{\mathrm{inv}}^{(d)})}$ $=1-\frac{4d^{3}+6nd+3d^{2}-d}{4d^{3}+6d^{2}+2d}=\frac{4d^{2}+6n+3d-1}{4d^{2}+6d+2}$
$=1-\frac{4d^{2}}{4d^{2}+6d+2}-\frac{3d+1}{4d^{2}+6d+2}-\frac{6n}{4d^{2}+6d+2}$
$=\Theta\left(\frac{1}{d}\right)-\frac{6n}{4d^{2}+6d+2}.$
Apparently, $6n/(4d^{2}+6d+2)$ always dominates $1/d$. In conclusion,
$1-\frac{\mathrm{Var}(X_{\mathrm{inv}}^{(d)})}{\mathrm{Var}(\hat{X}_{\mathrm{inv}}^{(d)})}=\Theta\left(\frac{n}{d^{2}}\right),$
giving the condition of $k_{n}\log(k_{n})=o(d^{2}/n)$ according to the arguments in the proof of [6, Theorem 4.1]. The remaining steps of the proof are the same.
In the long case, we write $\delta:=n-d$ and note that
$\mathrm{Var}(\hat{X}_{\mathrm{inv}})$ $=-\frac{1}{6}(n-\delta)^{3}+\frac{1}{3}(n-\delta)^{2}n-\frac{5}{36}n^{3}+\frac{16}{36}n^{2}\delta+O(n^{2}),$
$\mathrm{Var}(X_{\mathrm{inv}})$ $=-\frac{1}{6}(n-\delta)^{3}+\frac{1}{3}(n-\delta)^{2}n-\frac{5}{36}n^{3}+\frac{6}{36}n^{2}\delta+O(n^{2}).$
Therefore,
$1-\frac{\mathrm{Var}(X_{\mathrm{inv}})}{\mathrm{Var}(\hat{X}_{\mathrm{inv}})}=\frac{10n^{2}\delta+O(n^{2})}{\Theta(n^{3})}=\Theta\left(\max\left\{\frac{\delta}{n},\frac{1}{n}\right\}\right),$
giving the requirement of
$k_{n}\log(k_{n})\Theta\left(\max\left\{\frac{\delta}{n},\frac{1}{n}\right\}\right)=o(1)\Longleftrightarrow k_{n}\log(k_{n})=o\left(\min\left\{n,\frac{n}{\delta}\right\}\right).$
Again, the proof now follows the same steps as in [6, Theorem 4.1]. ∎
Remark 3.9.
This result is only useful if $d\gg n^{3/4}$. Otherwise, we can obtain the EVLT with a bound of $k_{n}=o(n^{1/2})$ from the universal EVLT [7, Theorem 5.1].
Extreme values of the joint distribution of $X_{\mathrm{inv}}^{(d)}$ and $X_{\mathrm{des}}^{(d)}$
The bivariate EVLT for $\left(X_{\mathrm{inv}}^{(d_{1})},X_{\mathrm{des}}^{(d_{2})}\right)^{\top}$ follows the same way, since the descent component does not interfere with any of the previously used arguments.
Theorem 3.10.
Let $(X_{n1},\ldots,X_{nk_{n}})$ be a row-wise i.i.d. triangular array with $X_{n1}\overset{\mathcal{D}}{=}\left(X_{\mathrm{inv}}^{(d_{1})},X_{\mathrm{des}}^{(d_{2})}\right)^{\top}$ on $S_{n}.$
  1. (a)
    If $d_{1}=o(n^{1/3}),$ then assume $k_{n}=\exp\bigl{(}o(n^{1/7}/d^{3/7})\bigr{)}.$
  2. (b)
    If $d_{1}$ satisfies the conditions in either Lemma 3.3 or 3.4, then assume $k_{n}\log(k_{n})=o(n^{-1}d_{1}^{2})$ in the short case and $k_{n}\log(k_{n})=o\left(\min\left\{n,\displaystyle{\frac{n}{n-d_{1}}}\right\}\right)$ in the long case.
Let $M_{n}:=\max\{X_{n1},\ldots,X_{nk_{n}}\}$ be the row-wise maximum. Let $\mu_{n}:=\mathbb{E}(X_{n1})$ and let $s_{n}:=\bigl{(}\sigma(X_{\mathrm{inv}}^{(d)}),\sigma(X_{\mathrm{des}})\bigr{)}$. Let $a_{n}:=s_{n}*\boldsymbol{\upalpha}_{k_{n}}$ and let $b_{n}:=s_{n}*\boldsymbol{\upbeta}_{k_{n}}+\mu_{n}$. Then,
$\forall\mathbf{x}\in\mathbb{R}^{2}\negthickspace:\quad\mathbb{P}(M_{n}\leq a_{n}*\mathbf{x}+b_{n})\longrightarrow\Lambda_{2}(\mathbf{x})\,.$
4. Proofs of Lemmas 3.3 and 3.4
We first consider $\hat{X}_{\mathrm{inv}}^{(d)}$ on symmetric groups and then provide the analogous observations on the other classical Weyl groups.
Proof of Lemma 3.3 for symmetric groups.
By (2), we have
$\mathbb{E}(X_{\mathrm{inv}}^{(d)}\mid Z_{k})=\sum_{(i,j)\in\mathfrak{N}_{n,d}}\mathbb{P}(Z_{i}>Z_{j}\mid Z_{k})=\sum_{(i,j)\in\mathfrak{N}_{n,d}}\begin{cases}1/2,&k\notin\{i,j\}\\ Z_{k},&k=i\\ 1-Z_{k},&k=j\end{cases}.$
Only the pairs $(i,j)$ with $k\in\{i,j\}$ contribute to $\mathrm{Var}\left(\mathbb{E}\left(X_{\mathrm{inv}}^{(d)}\mid Z_{k}\right)\right)$. These contributions are called the non-trivial parts for simplicity. The number of these pairs depends on whether $k$ belongs to $K_{1},K_{2},$ or $K_{3}$, as already seen in Remark 2.2. Figure 1 visualizes this case distinction for the exemplary choice of $n=15$ and $d=4$.
Figure 1. Overview of relevant pairs $(i,j)$ for computing the variance of $\mathbb{E}\left(X_{\mathrm{inv}}^{(d)}\mid Z_{k}\right)$, where $n=15$ and $d=4$. For each of the regions $K_{1},K_{2},K_{3},$ an exemplary index $k$ is chosen, and the pairs that give a non-trivial contribution are highlighted in red for $k\in K_{1}$, in blue for $k\in K_{2}$, and in green for $k\in K_{3}$.
If $k\in K_{2},$ then the non-trivial parts are
$\underbrace{Z_{k}+Z_{k}+\ldots+Z_{k}}_{d~{}\text{times}}+\underbrace{(1-Z_{k})+(1-Z_{k})+\ldots+(1-Z_{k})}_{d~{}\text{times}}=d.$
This means $\mathbb{E}\left(X_{\mathrm{inv}}^{(d)}\mid Z_{k}\right)$ is constant due to cancellation, and vanishes when computing the variance. So, in the short case, $\mathrm{Var}\left(\mathbb{E}\left(X_{\mathrm{inv}}^{(d)}\mid Z_{k}\right)\right)$ originates only from $K_{1}$ and $K_{3}$. If $k\in K_{1},$ then $\mathbb{E}\left(X_{\mathrm{inv}}^{(d)}\mid Z_{k}\right)=(k-1)(1-Z_{k})+dZ_{k}+\text{const}=(d+1-k)Z_{k}+\text{const}$. If $k\in K_{3}$, then $\mathbb{E}\left(X_{\mathrm{inv}}^{(d)}\mid Z_{k}\right)=(n-k)Z_{k}+d(1-Z_{k})+\text{const}=(n-d-k)Z_{k}+\text{const}$. So, the overall representation of $\hat{X}_{\mathrm{inv}}^{(d)}$ in the short case is
$\hat{X}_{\mathrm{inv}}^{(d)}=\sum_{k=1}^{n}\omega_{d}(k)Z_{k}+\text{const}\,,\qquad\text{with}\quad\omega_{d}(k):=\begin{cases}d-k+1\,,&k\in K_{1}\\ 0,&k\in K_{2}\\ n-d-k\,,&k\in K_{3}\end{cases}.$
Therefore,
$\mathrm{Var}\left(\hat{X}_{\mathrm{inv}}^{(d)}\right)$ $=\mathrm{Var}\left(\sum_{k=n-d+1}^{n}(n-k-d)Z_{k}+\sum_{k=1}^{d}(d+1-k)Z_{k}\right)$
$=\sum_{k=n-d+1}^{n}\frac{1}{12}(n-k-d)^{2}+\sum_{k=1}^{d}\frac{1}{12}(d+1-k)^{2}$
$=\frac{1}{12}\sum_{k=1}^{d}k^{2}+\frac{1}{12}\sum_{k=1}^{d}(-k)^{2}=\frac{1}{6}\frac{d(d+1)(2d+1)}{6}$
$=\frac{1}{72}\left(4d^{3}+6d^{2}+2d\right).$
So, the leading term is always $724d^{3}/$ 72. In light of Theorem 2.3, according to which $\mathrm{Var}\left(X_{\mathrm{inv}}^{(d)}\right)$ contains the monomials $726nd/$ 72 and $4d^{3}/72,$ it must be ensured that $d^{3}\gg nd\Longleftrightarrow d\gg\sqrt{n}$. ∎
Proof of Lemma 3.4 for symmetric groups.
Now, we consider the long case. Due to $d>n/2,$ we now have $n-d<d$, and the regions $K_{1},K_{2},K_{3}$ are redefined according to Remark 2.2. For the non-trivial parts, we note that:
  • If $k\in K_{1}$ or $k\in K_{3},$ then the non-trivial parts are the same as in the short case.
  • If $k\in K_{2},$ then the non-trivial parts yield $\mathbb{E}\left(X_{\mathrm{inv}}^{(d)}\mid Z_{k}\right)=(n-2k+1)Z_{k}+\text{const}$.
So, we again obtain a representation in the way of
$\hat{X}_{\mathrm{inv}}^{(d)}=\sum_{k=1}^{n}\omega_{d}(k)Z_{k}+\text{const}\,,\qquad\text{with}\quad\omega_{d}(k):=\begin{cases}d-k+1,&k\in K_{1}\\ n-2k+1,&k\in K_{2}\\ n-d-k,&k\in K_{3}\end{cases}.$
From this, we state that
(6a) $\mathrm{Var}\left(\hat{X}_{\mathrm{inv}}^{(d)}\right)$ $=\frac{1}{12}\sum_{k=1}^{n-d}(d+1-k)^{2}$
(6b) $+\frac{1}{12}\sum_{k=n-d+1}^{d}(n-2k+1)^{2}$
(6c) $+\frac{1}{12}\sum_{k=d+1}^{n}(n-k-d)^{2}.$
By appropriate index shifting, we calculate
[…]
$=\frac{1}{36}(n-d)\bigl{(}14d^{2}+d(9-10n)+12n^{2}-3n+1\bigr{)},$
[…]
This gives the total result
$\mathrm{Var}\left(\hat{X}_{\mathrm{inv}}^{(d)}\right)=\frac{1}{36}\left(-6d^{3}+12d^{2}n-9d^{2}-16dn^{2}+12dn-3d+11n^{3}-3n^{2}+2n\right).$
In contrast, by Theorem 2.3,
$\mathrm{Var}\left(X_{\mathrm{inv}}^{(d)}\right)=\frac{1}{36}\left(-6d^{3}+\left(12n-\frac{21}{2}\right)d^{2}-\left(6n^{2}-15n+\frac{9}{2}\right)d+n^{3}-3n^{2}+2n\right).$
Since the long case implies $n/2<d<n$, all monomials of order 3 are leading terms. The monomials $-16dn^{2}+11n^{3}$ appearing in $\mathrm{Var}\left(\hat{X}_{\mathrm{inv}}^{(d)}\right)$ and $-6n^{2}d+n^{3}$ appearing in $\mathrm{Var}\left(X_{\mathrm{inv}}^{(d)}\right)$ do not match in general. Using $\delta:=n-d$, we can write
$\mathrm{Var}(\hat{X}_{\mathrm{inv}}^{(d)})$ $=-\frac{1}{6}(n-\delta)^{3}+\frac{1}{3}(n-\delta)^{2}n-\frac{5}{36}n^{3}+\frac{16}{36}n^{2}\delta+O(n^{2}),$
$\mathrm{Var}(X_{\mathrm{inv}}^{(d)})$ $=-\frac{1}{6}(n-\delta)^{3}+\frac{1}{3}(n-\delta)^{2}n-\frac{5}{36}n^{3}+\frac{6}{36}n^{2}\delta+O(n^{2}).$
This shows that we need $\delta=o(n),$ since otherwise, there are two different leading coefficients of $n^{3}$. ∎
We now derive the analogous statements for the other classical Weyl groups $B_{n}$ and $D_{n}$. It is sufficient to prove the lemmas 3.3 and 3.4 for the groups $B_{n},$ since the difference between $X_{\mathrm{inv}}^{B,(d)}$ and $X_{\mathrm{inv}}^{D,(d)}$ is asymptotically negligible (cf. (5a) and (5b)). Recall the asymptotic quantification of $\mathrm{Var}(X_{\mathrm{inv}}^{(d)})$ given in Lemma 2.6.
To compute $\mathrm{Var}(\hat{X}_{\mathrm{inv}}^{(d)}),$ we ignore all constant parts appearing in $\hat{X}_{\mathrm{inv}}^{(d)}$. By (5a), we can split ${\mathbb{E}(X_{\mathrm{inv}}^{B,(d)}\mid Z_{k})=\mathbb{E}(X_{\mathrm{inv}}^{B,(d)}\mid Z_{k})^{+}+\mathbb{E}(X_{\mathrm{inv}}^{B,(d)}\mid Z_{k})^{-}+\mathbb{E}(X_{\mathrm{inv}}^{B,(d)}\mid Z_{k})^{\circ},}$ where
$\mathbb{E}(X_{\mathrm{inv}}^{B,(d)}\mid Z_{k})^{+}$ $:=\sum_{(i,j)\in\mathfrak{N}_{n,d}}\mathbb{P}(Z_{i}>Z_{j}\mid Z_{k})\,,$
$\mathbb{E}(X_{\mathrm{inv}}^{B,(d)}\mid Z_{k})^{-}$ $:=\sum_{(i,j)\in\widetilde{\mathfrak{N}}_{n,d}}\mathbb{P}(-Z_{i}>Z_{j}\mid Z_{k})\,,$
$\mathbb{E}(X_{\mathrm{inv}}^{B,(d)}\mid Z_{k})^{\circ}$ $:=\sum_{i=1}^{n}\mathbb{P}(Z_{i}<0\mid Z_{k})\,.$
The third sum ${\mathbb{E}(X_{\mathrm{inv}}^{B,(d)}\mid Z_{k})^{\circ}}$ is asymptotically negligible. So, to obtain a representation $\hat{X}_{\mathrm{inv}}^{B,(d)}=\displaystyle{\sum\nolimits_{i=1}^{n}\omega_{d}(k)Z_{k}}+\text{const},$ we write ${\omega_{d}(k)=\omega_{d}(k)^{+}+\omega_{d}(k)^{-}}$, with ${\omega_{d}(k)^{+}}$ stemming from ${\mathbb{E}(X_{\mathrm{inv}}^{B,(d)}\mid Z_{k})^{+}}$ and ${\omega_{d}(k)^{-}}$ stemming from ${\mathbb{E}(X_{\mathrm{inv}}^{B,(d)}\mid Z_{k})^{-},}$ respectively.
For ${\omega_{d}(k)^{+}},$ we can use the previous counting method. However, we have to take into account that on $B_{n}$,
$\mathbb{P}(Z_{k}>Z_{j}\mid Z_{k})$ $=\frac{Z_{k}+1}{2}\,,$ $\mathbb{P}(Z_{i}>Z_{k}\mid Z_{k})$ $=\frac{1-Z_{k}}{2}\,,$
$\mathbb{P}(-Z_{k}>Z_{j}\mid Z_{k})$ $=\frac{1-Z_{k}}{2}\,,$ $\mathbb{P}(Z_{i}>Z_{k}\mid Z_{k})$ $=\frac{1-Z_{k}}{2}\,.$
In conclusion, the coefficients ${\omega_{d}(k)^{+}}$ on $B_{n}$ are half of the coefficients $\omega_{d}(k)$ on $S_{n}$. Note that ${\omega_{d}(k)^{+}}=(n-2k+1)/2$ for all $d\geq n$. Moreover, ${\omega_{d}(k)^{-}}=-\widetilde{N}_{n,d}^{(k)}/2,$ where
$\widetilde{N}_{n,d}^{(k)}:=|\{(i,j)\in\widetilde{\mathfrak{N}}_{n,d}\mid\text{either}~{}i=k~{}\text{or}~{}j=k\}|\,.$
Recall that by Remark 2.5, the short case is now $d<n$ and the long case is $d\geq n$. By counting the number of such $(k,j)$ and $(i,k)$, we can determine the asymptotic quantification of $\mathrm{Var}\left(\hat{X}_{\mathrm{inv}}^{B,(d)}\right)$.
Proof of Lemma 3.3 for the groups $B_{n}$ and $D_{n}$.
If $d\leq n/2$, then all pairs in $\widetilde{\mathfrak{N}}_{n,d}$ are located within $K_{1}$, yielding
$\omega_{d}(k)$ =$=\begin{cases}(d-k+1)/2-(d-k-1)/2,&k\leq d/2\\ (d-k+1)/2-(d-k)/2,&d/2<k\leq d\\ 0,&d<k\leq n-d\\ (n-d-k)/2,&n-d<k\leq n\end{cases}$
=$=\begin{cases}(n-d-k)/2,&n-d<k\leq n\\ O(1),&\text{otherwise}\end{cases}\,.$
In conclusion, if $d\leq n/2,$ then
$\mathrm{Var}\left(\hat{X}_{\mathrm{inv}}^{B,(d)}\right)$ $=\frac{1}{3}\sum_{k=1}^{n}\omega_{d}(k)^{2}+O(d^{2})=\frac{1}{12}\sum_{k=n-d+1}^{n}(n-d-k)^{2}+O(d^{2})$
$=\frac{1}{12}\frac{d(d+1)(2d+1)}{6}+O(d^{2})=\frac{1}{36}d^{3}+O(d^{2}).$
Due to $\mathrm{Var}\left(X_{\mathrm{inv}}^{B,(d)}\right)=d^{3}/36+nd/12+O(d^{2})$ according to Lemma 2.6, we obtain the same condition as in Lemma 3.3, namely, $d\gg\sqrt{n}$.
If $n/2<d<n$, then the pairs in $\widetilde{\mathfrak{N}}_{n,d}$ also cover $K_{2}$. For $k\in K_{2},$ there cannot be any pairs $(k,j)$ if $n/2<d\leq 2n/3,$ while this is possible if $d>2n/3$. However, the difference between these two subcases is only marginal. If $n/2<d\leq 2n/3,$ we obtain
$\omega_{d}(k)=\begin{cases}O(1),&k\in K_{1}\\ \bigl{(}n-2k+1-(d-k)\bigr{)}/2,&k\in K_{2}\\ (n-d-k)/2,&k\in K_{3}\end{cases}\,.$
If $d>2n/3,$ then
$\omega_{d}(k)=\begin{cases}O(1),&k\in K_{1}\\ \bigl{(}n-2k+1-(d-k-1)\bigr{)}/2,&k\in K_{2},k\leq d/2\\ \bigl{(}n-2k+1-(d-k)\bigr{)}/2,&k\in K_{2},k>d/2\\ (n-d-k)/2,&k\in K_{3}\end{cases}\,.$
In conclusion, if $n/2<d<n,$ then
$\mathrm{Var}\left(\hat{X}_{\mathrm{inv}}^{B,(d)}\right)$ $=\frac{1}{3}\sum_{k=1}^{n}\omega_{d}(k)^{2}+O(d^{2})$
$=\frac{1}{12}\sum_{k=n-d+1}^{d}(n-d-k+1)^{2}+\frac{1}{12}\sum_{k=d+1}^{n}(n-d-k)^{2}+O(d^{2})$
$=\frac{1}{12}\sum_{k=n-d+1}^{d}(n-d-k+1)^{2}+\frac{1}{12}\sum_{k=1}^{2d-n}k^{2}+O(d^{2})$
$=\frac{1}{12}\sum_{k=1}^{2d-n-1}k^{2}+\frac{1}{12}\left(\sum_{k=1}^{d}k^{2}-\sum_{k=1}^{2d-n}k^{2}\right)+O(d^{2})$
$=\frac{1}{36}d^{3}+O(d^{2})\,.\qed$
Proof of Lemma 3.4 for the groups $B_{n}$ and $D_{n}$.
In the long case, the main focus is on counting $\widetilde{N}_{n,d}^{(k)}$. Figure 2 illustrates the positions of pairs $(i,j)\in\widetilde{\mathfrak{N}}_{n,d}$ in the long case for the exemplary choice of $n=12,d=16$.
Figure 2. Visualization of pairs $(i,j)$ in $\widetilde{\mathfrak{N}}_{n,d}$ in the long case for $n=12$ and $d=16$. The numbers $d-n$ and $d/2$ are important case distinction thresholds for counting the pairs $(k,j)$ and $(i,k),$ respectively.
With help of Figure 2, it is straightforward to count
$\widetilde{N}_{n,d}^{(k)}=\begin{cases}n-1,&1\leq k\leq d-n\\ d-k-1,&d-n<k\leq d/2\\ d-k,&d/2<k\leq n\end{cases}\,.$
This result is also illustrated in Figure 3, which displays the number of pairs $(i,k)$ and the number of pairs $(k,j)$.
Figure 3. Plots of the numbers of pairs $(i,k)$ (red) and $(k,j)$ (blue) in $\widetilde{\mathfrak{N}}_{n,d}$. The sum of these two numbers is $\widetilde{N}_{n,d}^{(k)},$ which is displayed by the black crossed points.
Therefore, in the long case, we have
$\omega_{d}(k)=\begin{cases}1-k,&1\leq k\leq d-n\\ (n+2-d-k)/2,&d-n<k\leq d/2,\\ (n+1-d-k)/2,&d/2<k\leq n.\end{cases}$
We compute
$\sum_{k=1}^{n}\omega_{d}(k)^{2}$ $=\sum_{k=1}^{d-n}(k-1)^{2}+\frac{1}{4}\sum_{k=d-n+1}^{n}(n-d+1-k)^{2}+O(n^{2})$
$=\sum_{k=1}^{d-n-1}k^{2}+\frac{1}{4}\sum_{k=d-n+1}^{n}\bigl{(}k-(d-n-1)\bigr{)}^{2}+O(n^{2})$
$=\frac{1}{3}(d-n-1)^{3}+\frac{1}{12}(2n-d-1)^{3}+O(n^{2})$
$=\frac{d^{3}}{4}-\frac{d^{2}n}{2}-\frac{3d^{2}}{4}+dn+\frac{3d}{4}+\frac{n^{3}}{3}-\frac{n}{2}+\frac{1}{4}+O(n^{2})\,.$
Due to $\mathrm{Var}(Z_{k})=1/3,$ we obtain
$\mathrm{Var}(\hat{X}_{\mathrm{inv}}^{B,(d)})=\frac{d^{3}}{12}-\frac{d^{2}n}{6}+\frac{n^{3}}{9}+O(d^{2}).$
On the other hand, by Lemma 2.6,
$\mathrm{Var}(X_{\mathrm{inv}}^{B,(d)})=-\frac{1}{12}d^{3}+\frac{1}{3}nd^{2}-\frac{1}{3}n^{2}d+\frac{1}{9}n^{3}+O(d^{2})\,,$
i.e., the leading terms $d^{3}/12-d^{2}n/6$ and $-d^{3}/12+nd^{2}/3-n^{2}d/3$ need to match. Writing $d=cn(1+o(1))$ for $1\leq c\leq 2,$ this gives the condition
$\frac{(cn)^{3}}{12}-\frac{c^{2}n^{3}}{6}=-\frac{(cn)^{3}}{12}+\frac{c^{2}n^{3}}{3}-\frac{cn^{3}}{3}\Longleftrightarrow\frac{c^{3}}{6}-\frac{c^{2}}{2}+\frac{c}{3}=0\,,$
which holds precisely for $c=1$ and $c=2$. ∎